A 2.58 g lead weight, initially at 10.4 ∘C, is submerged in 7.85 g of water at 52.3 ∘C in an insulated container. e What is the final temperature of both substances at thermal equilibrium?
Answer #1

m(lead) = 2.58 g

T(lead) = 10.4 oC

C(lead) = 0.128 J/goC

m(water) = 7.85 g

T(water) = 52.3 oC

C(water) = 4.184 J/goC

T = to be calculated

Let the final temperature be T oC

use:

heat lost by water = heat gained by lead

m(water)*C(water)*(T(water)-T) = m(lead)*C(lead)*(T-T(lead))

7.85*4.184*(52.3-T) = 2.58*0.128*(T-10.4)

32.8444*(52.3-T) = 0.3302*(T-10.4)

1717.7621 - 32.8444*T = 0.3302*T - 3.4345

T= 51.8829 oC

Answer: 51.88 oC